OpenStudy (anonymous):

Choose the correct procedure for calculating (f − g)(5) if f(x) = x2 − 13 and g(x) = x + 5. show steps or explain please

OpenStudy (anonymous):

\[\left( f-g \right)\left( x \right)=f \left( x \right)-g \left( x \right)\] \[\left( f-g \right)\left( 5 \right)=f \left( 5 \right)-g \left( 5 \right)\] i think now you can solve

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

\[f \left( 5 \right)=5^{2} -13=25-13=12\] g(5)=5+5=10 (f-g)5=12-10=2